1 attempts left Check my work Enter your answer in the provided box. Ammonia (NH3) decomposes to hydrogen and nitrogen and 22.0 kcal/mol of energy is absorbed. 2 NHj(g) 3 H2(8) + N2(g) AH +22.0 kcal/mol How much energy is absorbed when 2.63 g of NHj reacts? keal of energy absorbed Report your answer to appropriate number of significant figures.
Answer #1

Molar mass of NH3,

MM = 1*MM(N) + 3*MM(H)

= 1*14.01 + 3*1.008

= 17.034 g/mol

mass(NH3)= 2.63 g

use:

number of mol of NH3,

n = mass of NH3/molar mass of NH3

=(2.63 g)/(17.03 g/mol)

= 0.1544 mol

Since Δ H is positive, heat is absorbed

when 2 mol of NH3 reacts, heat absorbed = 22.0 Kcal

So,

for 0.1544 mol of NH3, heat absorbed = 0.1544*22.0/2 Kcal

= 1.698 Kcal

Answer: 1.70 Kcal